# Data-Model Co-Evolution: Growing Test Sets to Refine LLM Behavior

### v1

@@ 5.1. Study Design
We employed a within-subjects design, with each participant completing a task in two conditions: our Co-Evolution workflow and a Baseline interface. To mitigate potential learning and carryover effects, we used two distinct yet comparable task domains: content moderation for campus-community comments and for restaurant-review comments. The order of conditions and the assignment of domains to each condition were counterbalanced across participants using a balanced Latin square. To further reduce domain-specific priming, participants completed a tutorial with a third, distinct domain (YouTube comments) before starting each of the two experimental blocks.
@@ 5.2. Task: Refining a Content Moderation Policy
The Co-Evolution workflow condition provided participants with our full system. This included features for generating synthetic test examples, exploring semantically similar examples around a point of failure, and refining the prompt instruction based on suggestions from the system.
The Baseline condition was a controlled, ablated version of our system that represented a standard prompt-editing environment. The user interface is looking similar, but with its data-centric workflow disabled. Specifically, it included the prompt editing and history panel on the left and the test set panel on the top right side of the interface where participants could manually add potential user inputs and observe model outputs on these manually crafted inputs. However, the baseline did not offer tools to generate test cases with LLMs, explore the neighborhood space around a failure, or use AI-assisted prompt suggestions.
@@ 5.4. Participants
We recruited 12 participants (3 female; mean age $25.7$, SD $2.4$) from mailing lists for university students in South Korea. Participants were required to have used a generative AI service (e.g., ChatGPT) at least ten times in the preceding week and to have completed at least one university-level AI/ML course or have equivalent experience. These criteria ensured participants were familiar with prompt-based interaction and possessed a foundational understanding of AI concepts, reflecting our target users who are not necessarily AI experts but have sufficient background to engage in prompt refinement and test set construction. Each participant received a gift card (KRW 20,000 $\approx$ USD 15) for their time.
@@ 5.5. Session Protocol
Each session lasted approximately 55 minutes and was conducted remotely via video conferencing. After providing electronic consent and completing a brief pre-study questionnaire, participants proceeded through the two counterbalanced experimental blocks. Each block began with a short tutorial and Q&A session on the assigned condition. Participants then completed the task for up to 10 minutes. Following the task, they spend approximately 3 minutes labeling a holdout evaluation set (detailed in the next paragraph). Then they filled out a block-specific post-questionnaire consisting of 10 questions with 7-point Likert scales and 3 open-ended questions.
@@ 6.1. RQ 1: How do prompt instructions refined by our workflow differ from the baseline?
Figure 5. Length of prompt instructions created under each condition, measured in word count and sentence count. Each point represents one participant (jittered for visibility), with box plots summarizing distributions. Instructions produced with the Co-Evolution workflow were significantly longer than those from the Baseline.
@@
Prompt instructions created with the Co-Evolution workflow were significantly longer than those from the Baseline. The average word count was 112.4, compared to 82.2 in the Baseline condition. A paired $t$-test confirmed this difference was statistically significant ($p<0.001$). This trend was consistent across both task domains (Campus: 112.5 vs. 79.5 words; Reviews: 112.3 vs. 84.8 words). Sentence counts were also higher under Co-Evolution (10.0 vs 7.9). Because a Shapiro-Wilk test (Shapiro and Wilk, 1965) indicated borderline normality ($p=0.045$), we ran a Wilcoxon signed-rank test, which again confirmed the difference ($p=0.031$).
@@
To measure specificity, we assessed the following three structural features of the prompt instructions: (1) the number of policy statements, i.e., individual sentences articulating specific rules or criteria about the model behavior; (2) the presence of explicit exceptions or constraints, such as “do not” rule or boundaries of the task; and (3) the inclusion of concrete examples that illustrate policies.
The results are shown in Figure 6. The average number of policy statements did not differ significantly between the two conditions (i.e., 2.83 for Baseline vs 3.25 for Co-Evolution; paired $t$-test: $p=0.435$). However, prompt instructions from the Co-Evolution workflow more frequently included explicit exceptions or constraints (9 vs. 6 out of 12 participants) and concrete examples (8 vs. only 1). These results suggest that our workflow helps participants articulate boundaries and illustrate rules with examples, producing instructions that are not only longer but also richer in detail.
@@ 6.2. RQ 2: How does the workflow shape the prompt refinement process?
Participants produced a comparable number of revisions in the two conditions. On average, they saved 3.92 versions in Baseline and 4.08 versions in Co-Evolution, a difference that was not statistically significant (distributions shown in Figure 7).
Figure 7. Prompt refinement process measured by the number of instruction versions (left) and the test set size (right). Participants produced a similar number of versions across conditions, and those using Co-Evolution generated substantially larger test sets than in the Baseline.
However, a significant difference was observed in the size of the test sets created, as expected thanks to the neighborhood probing feature. Participants using the Co-Evolution workflow generated substantially larger test sets (avg. 13.42 items) compared to those in the Baseline (avg. 7.17 items). A Wilcoxon signed-rank test confirmed this difference was highly significant ($p<0.001$).
@@ 6.3. RQ 3: Do instruction revisions lead to changes in model behavior?
In the Baseline condition, participants produced an average of 3.0 outcome changes relative to the default prompt instruction. Of these, 47% shifted to “not problematic.” In the Co-Evolution condition, the average was higher at 3.9 changes with 45% moving to “not problematic.” The difference in counts was not statistically significant.
@@
The distribution of changes, however, revealed important contrasts (shown in Figure 8). In the Baseline condition, four participants produced no changes at all, suggesting that despite effort, their final instructions had little measurable impact on the model behavior. Among those who did produce changes, the most active participants tended to alter outcomes in a highly one-sided fashion (e.g., all six changes are from “not problematic” to “problematic”). By contrast, in the Co-Evolution condition, every participant induced at least one change, and the distribution was less polarized. This suggests that while both interfaces can alter model behavior, Co-Evolution led to more consistent and balanced effects across participants.
@@ 6.4. RQ 4: Does the workflow improve alignment between instructions and model behavior?
Figure 9. Alignment between model outputs and participant labels, measured by F1-score. Participants’ revised instructions in Co-Evolution yielded higher alignment than in the Baseline. A paired $t$-test indicated this difference was marginally significant at the 90% confidence level.
@@
While the Baseline condition achieved an average F1-score of 0.58, our Co-Evolution workflow achieved 0.68 (distributions shown in Figure 9). To assess whether this improvement was statistically reliable, we first conducted a Shapiro-Wilk test of normality, which indicated no violation ($p=0.43$). A paired $t$-test then showed that the improvement was marginally significant at the 90% confidence level ($p=0.0879$).
@@
These results highlight that the instruction revisions observed in RQ 3 were not arbitrary: they shifted the model’s behavior closer to participants’ intended judgments. Although the improvement was only marginally significant, the findings suggest that the co-evolution workflow can guide participants toward revisions that both alter model behavior and bring it into better alignment with their goals.
6.5. RQ 5: How do participants perceive the effectiveness of the workflow?
@@
The above four research questions analyzed what changed in the prompt instructions and how those changes affected model behavior. We now shift from these outcome-focused measures to participants’ own subjective assessments.
@@
Figure 10. Subjective ratings for 10 questionnaire items, comparing the Baseline and Co-evolution conditions. Bars represent the distribution of responses on a 7-point Likert scale, with darker shades indicating stronger agreement. Average scores are shown to the right of each bar, and significance markers denote differences between conditions, with ** indicating $p<.01$ and * indicating $p<.1$.
@@
Figure 10 shows the distribution of subjective ratings for the 10 questions for each of the two conditions. First, both conditions scored highly on basic usability. The baseline, with fewer features, was rated slightly higher on ease of learning, as expected (Q1), and both conditions were judged equally easy to use with a high average score (Q2; mean 6.08 for both). Even on sense of control (Q3), where we anticipated the baseline might be stronger, ratings were comparable, with co-evolution even a bit higher (mean: 6.08 vs. 5.25). These patterns suggest that the two systems were broadly on par for straightforward ease of use.
@@
Where the co-evolution workflow stood out was in the dimensions it was designed to support. Participants reported that it helped adapt the system to their own preferences and criteria, with 10 of 12 gave ratings of 6 (agree) or 7 (strongly agree) compared to only 2 under the baseline (Q4; $p<0.01$). They also reported greater confidence that the AI would behave as intended (Q5; $p<0.01$) and found the workflow more effective for covering the “long tail”, such as discovering unseen cases (Q6; $p<0.1$) and defining policy boundaries in realistic scenarios (Q7; $p<0.01$). Participants also emphasized that the workflow clarified their policy decisions where 9 vs. 5 participants rated at 6 or 7 (Q8; $p<0.01$) and left them more satisfied overall, with very stronger intentions to reuse it for future tasks (Q10; $p<0.01$).
@@ 7. Discussion
Our user study provides valuable insights into how our interactive systems can scaffold the complex, iterative process of data-model co-evolution. Based on our qualitative observations and participant feedback, we discuss how our system facilitates this process and reflect on its limitations, which in turn suggest critical design implications for future systems.
@@ 7.1.1. Enabling Fine-Grained Specification through Concrete Examples
A significant advantage reported by participants was the ability to define nuanced, domain-specific policies by directly interacting with concrete data examples. This was particularly crucial for ambiguous tasks like content moderation, where abstract guidelines often fail to capture real-world complexities. By reviewing and labeling diverse examples, participants could more easily identify and articulate the subtle boundaries of their criteria. One participant noted how this process helped them refine their standards: “By creating a diverse test set for ‘Bad’ items where the boundary was ambiguous, I was able to consider more detailed criteria.”
This hands-on approach, grounded in tangible data, transformed the abstract task of policy-making into a more manageable and effective process. Another participant confirmed this, stating: “Seeing diverse expressions [of criticism] that I could encounter in real life with my own eyes was very helpful in setting more detailed standards for the acceptable level of criticism.”
@@ 7.1.2. Facilitating Nuanced and Rapid Co-evolution
Participants highly valued the system’s interactivity. The rapid feedback loop allowed them to quickly test hypotheses and iterate on their prompts, making the refinement process more efficient and intuitive. One participant remarked: “And it was impressive that it processed feedback in real-time and let me modify the prompt command again.”
The comparison table view was essential for helping users understand the concrete impact of their changes. By visualizing the outputs of the old and new prompt instructions simultaneously, users could easily assess whether their modifications led to the desired improvements. This feature directly supported the goal-oriented nature of their task. “Being able to compare the results of the old prompt and the new prompt at a glance was helpful in writing the prompt’s content (e.g., community guidelines) because I could see what difference the new prompt made.” This underscores the importance of not just enabling rapid iteration but also providing tools that make the outcomes of those iterations understandable.

### v2

@@ 6.1. Study Design
We employed a within-subjects design, with each participant completing a task in two conditions: our Co-Evolution workflow and a baseline interface. To mitigate potential learning and carryover effects, we used two distinct yet comparable task domains: content moderation for campus community websites and for restaurant reviews. The order of conditions and the assignment of domains to each condition were counterbalanced across participants using a balanced Latin square. To further reduce domain-specific priming, participants completed a tutorial with a third, distinct domain (YouTube comments) before starting each of the two experimental blocks. The study was approved by our institution’s IRB.
@@ 6.2. Task: Refining a Content Moderation Policy
The Co-Evolution workflow condition provided participants with our full system. This included features for generating synthetic test examples, exploring similar examples around a point of failure, and refining prompt instructions based on suggestions from the system.
The baseline condition was a controlled, ablated version of our system that represented a standard prompt-editing environment. The user interface is similar, but with its data-centric workflow disabled. Specifically, it included the prompt editing and history panel on the left and the test set panel on the top right side of the interface where participants could manually add potential user inputs and observe model outputs on these manually crafted inputs. However, the baseline did not offer tools to generate test cases with LLMs, explore the neighborhood space around a failure, or use AI-assisted prompt suggestions.
To ensure a fair comparison and mitigate the difficulty of initiating the test set creation process (“cold start”), participants in both conditions were initially provided with an identical set of three pre-populated test examples. This provided all participants with a consistent initial anchor for their refinement process.
@@ 6.4. Participants
We recruited 16 participants (5 female; mean age 25.3, SD 2.3) from mailing lists for university students in South Korea. We set two inclusion criteria: participants were required to have used a generative AI service (e.g., ChatGPT) at least ten times in the preceding week; and to have completed at least one university-level AI/ML course or have equivalent experience. These criteria ensured participants were familiar with prompt-based interaction and possessed a foundational understanding of AI concepts, reflecting our target users who are not necessarily AI experts but have sufficient background to engage in prompt refinement and test set construction. We deliberately targeted this demographic to reflect the evolving definition of developers or practitioners. As LLMs make application development more accessible, developers are increasingly individuals who are not necessarily professional software engineers but understand core AI concepts, specifically test sets and evaluation. All participants were familiar with Korean, and the pre-populated prompt instructions and test examples were provided in Korean. Each participant received a gift card (KRW 20,000 $\approx$ USD 15) as compensation.
@@ 6.5. Session Protocol
Each session lasted approximately 55 minutes and was conducted remotely via video conferencing. After providing electronic consent and completing a brief pre-study questionnaire, participants proceeded through the two counterbalanced experimental blocks. Each block began with a short tutorial and Q&A session. Participants then completed the task for up to 10 minutes. Following the task, they spent approximately 3 minutes labeling a holdout evaluation set (detailed in the next paragraph). Then they filled out a block-specific post-questionnaire consisting of 10 questions with 7-point Likert scales and 3 open-ended questions.
Figure 5. Length of prompt instructions created under each condition, measured in word count and sentence count. Each point represents one participant (jittered for visibility), with box plots summarizing distributions. Instructions produced with the Co-Evolution workflow were significantly longer than those from the baseline condition.
The figure includes two charts, one for word count and one for sentence count. Each chart displays two horizontal box plots, with the baseline condition above and the Co-Evolution condition below, and 16 jittered data points per condition. In both panels, values under the Co-Evolution condition are higher than those under the baseline.
The figure contains three charts: (1) Number of Policy Statements: this chart shows horizontal box plots, with the baseline condition above and the Co-Evolution condition below. The two conditions have similar medians and upper quartiles, with modest differences in the lower quartile and maximum values. (2) Presence of Explicit Exceptions or Constraints: this chart shows a horizontal bar chart with one bar per condition, indicating that the measured value is slightly higher for the Co-Evolution condition. (3) Inclusion of Concrete Examples: this chart shows a horizontal bar chart with one bar per condition, indicating a substantially higher measured value under the Co-Evolution condition.
@@ 6.6. Holdout Evaluation Sets
Figure 7. Prompt refinement process measured by the number of instruction versions (left) and the test set size (right). Participants produced a similar number of versions across conditions, and those using Co-Evolution generated substantially larger test sets than in the baseline condition.
The figure includes two box-plot panels comparing conditions. One panel shows the number of prompt instruction versions, and the other shows the number of test set examples. In both panels, each condition is displayed as a horizontal box plot, with the baseline condition above and the Co-Evolution condition below. For prompt instruction versions, the two conditions show similar distributions, with similar medians and interquartile ranges. For test set examples, the Co-Evolution condition shows higher values than the baseline.
@@ 7.1. RQ 1: How do prompt instructions refined by our workflow differ from the baseline?
Prompt instructions created with the Co-Evolution workflow were significantly longer than those from the baseline. The average word count was 111.1, compared to 81.2 in the baseline condition. A paired $t$-test confirmed this difference was statistically significant ($p<0.001$). This trend was consistent across both task domains (campus: 107.8 vs. 78.9 words; reviews: 114.5 vs. 83.5 words). Sentence counts were also higher under Co-Evolution (9.8 vs 7.9), and a paired $t$-test indicated that the difference is significant ($p=0.019$).
@@
To measure specificity, we assessed three structural features of the prompt instructions: (1) the number of policy statements, i.e., individual sentences articulating specific rules or criteria about the model behavior; (2) the presence of explicit exceptions or constraints, such as “do not” rule or boundaries of the task; and (3) the inclusion of concrete examples that illustrate policies.
The results are shown in Figure 6. The average number of policy statements did not differ significantly between the two conditions (i.e., 3.00 for the baseline vs. 3.31 for Co-Evolution). However, prompt instructions from the Co-Evolution workflow more frequently included explicit exceptions or constraints (13 vs. 9 out of 16 participants) and concrete examples (9 vs. only 2). These results suggest that our workflow helps participants articulate boundaries and illustrate rules with examples, producing instructions that are not only longer but also richer in detail.
@@ 7.2. RQ 2: How does the workflow shape the prompt refinement process?
Participants produced a comparable number of revisions in the two conditions. On average, they saved 3.88 versions in the baseline and 4.31 versions in Co-Evolution, a difference that was not statistically significant (distributions shown in Figure 7).
However, a significant difference was observed in the size of the test sets created, as expected thanks to the neighborhood probing feature. Participants using the Co-Evolution workflow generated substantially larger test sets (avg. 14.19 items) compared to those in the baseline (avg. 7.94 items). A $t$-test confirmed this difference was highly significant ($p<0.001$).
@@ 7.3. RQ 3: Do instruction revisions lead to changes in model behavior?
In the baseline condition, participants produced an average of 3.25 outcome changes relative to the default prompt instruction. In the Co-Evolution condition, the average was higher at 4.25 changes. The difference in counts was not statistically significant. The shape of the distribution, however, revealed contrasts (shown in Figure 8). In the baseline condition, four participants produced no changes at all, and the Shapiro-Wilk test indicated a non-normal distribution. In contrast, in the Co-Evolution condition, every participant induced at least one change, and the distribution satisfied normality.
@@
This chart shows paired horizontal box plots comparing the number of changes across conditions. The baseline condition includes four observations at zero, with the remaining values extending up to six. In contrast, the Co-Evolution condition exhibits a more evenly distributed range of values, spanning from one to nine.
@@ 7.4. RQ 4: Does the workflow improve alignment between instructions and model behavior?
While the baseline condition achieved an average F1-score of 0.56, our Co-Evolution workflow achieved 0.69 (distributions shown in Figure 9). To assess whether this improvement was statistically reliable, we first conducted a Shapiro-Wilk test of normality, which indicated no violation ($p=0.33$). A paired $t$-test then showed that the improvement was significant at the 95% confidence level ($p=0.0241$).
@@
These results highlight that the instruction revisions observed in RQ 3 were not arbitrary: they shifted the model’s behavior closer to participants’ intended judgments. They suggest that the Co-Evolution workflow can guide participants toward revisions that both alter model behavior and bring it into better alignment with their goals.
To rule out the possibility that these results were influenced by task-order learning effects, we performed additional analyses. First, to examine whether participants improved simply by performing one task before the other, for each participant, we computed the accuracy difference between the second and first task. Performance on the first task was slightly higher on average, and the difference was not statistically significant ($p=0.31$). We also compared the advantage of the Co-Evolution condition across interface orderings (participants who began with the baseline vs. those who began with Co-Evolution). Again, we found no significant differences (Welch’s $t$-test $p=0.74$), indicating that the observed benefit of Co-Evolution is unlikely to result from learning effects.
Figure 9. Alignment between model outputs and participant labels, measured by F1-score. Participants’ revised instructions in Co-Evolution yielded higher alignment than in the baseline. A paired $t$-test indicated this difference was significant at the 95% confidence level ($p=0.0241$).
This chart shows paired horizontal box plots comparing F1-scores across conditions. The Co-Evolution condition has a higher median and higher first and third quartiles than the baseline. While the maximum F1-score is slightly higher in the baseline condition, the Co-Evolution condition shows a narrower overall range, with a substantially higher minimum value.
7.5. RQ 5: What are the efficacy and coverage of the test sets created?
To investigate the quality of the constructed test sets, we analyzed the data along two dimensions: efficacy (measuring how effectively examples exposed model limitations) and coverage (measuring the semantic diversity of the test scenarios).
First, to assess the efficacy of the test sets, we examined whether the generated examples actively probed model limitations rather than being trivial inputs. We measured the rate at which test inputs successfully triggered a model failure—specifically, examples participants explicitly marked as incorrect at least once during the session. The Co-Evolution workflow demonstrated a significant advantage: 51.9% of the test examples (93 out of 179) were flagged as failures, whereas the baseline condition yielded only 26.7% (23 out of 86). This indicates that the Co-Evolution workflow more effectively guides users to generate challenging edge cases that probe the model’s weaknesses.
Second, to determine the coverage of the test sets, we examined the semantic diversity of inputs to ensure they addressed a broad range of scenarios. Specifically, for each participant, we calculated the coverage based on the four categories (e.g., target group, problematic style) defined in our technical experiment (Section 5). The Co-Evolution condition achieved a significant higher coverage of 58% (avg. 13.8 out of 24 categories), compared to the baseline, which covered 49% (avg. 11.7 categories). While this increase is partly attributable to the higher total number of examples generated in the Co-Evolution condition, the result suggests that the workflow successfully scaffolded participants in broadening their data space to cover a more diverse range of scenarios, rather than simply generating more of the same types of errors.
7.6. RQ 6: How do participants perceive the effectiveness of the workflow?
@@
The above five research questions analyzed what changed in the prompt instructions and how those changes affected model behavior. We now shift from these outcome-focused measures to participants’ own subjective assessments.
@@
Figure 10. Subjective ratings for 10 questionnaire items, comparing the baseline and Co-Evolution (Co-Evo) conditions. Bars represent the distribution of responses on a 7-point Likert scale, with darker shades indicating stronger agreement. Average scores are shown to the right of each bar, and significance markers denote differences between conditions, with * indicating $p<0.05$ and ** indicating $p<0.01$.
This figure shows ten questionnaire items presented as paired horizontal percentage stacked bar charts, with the baseline condition above and the Co-Evolution condition below. Each bar represents the distribution of responses on a 7-point Likert scale. Mean scores are displayed to the right of each chart, and asterisks indicate statistically significant differences between the conditions.
@@
Figure 10 shows the distribution of subjective ratings for the 10 questions for each of the two conditions. First, both conditions scored highly on basic usability. The baseline, with fewer features, was rated slightly higher on ease of learning, as expected (Q1), and both conditions were judged similarly easy to use (Q2; 5.94 vs. 6.19). Even on sense of control (Q3), where we anticipated the baseline might be stronger, ratings were comparable, with Co-Evolution even a bit higher (mean: 5.94 vs. 5.31). These patterns suggest that the two systems were broadly on par for straightforward ease of use.
@@
Where the Co-Evolution workflow stood out was in the dimensions it was designed to support. Participants reported that it helped adapt the system to their own preferences and criteria, with 13 of 16 gave ratings of 6 (agree) or 7 (strongly agree) compared to 4 under the baseline (Q4; $p<0.01$). They also reported greater confidence that the AI would behave as intended (Q5; $p<0.05$) and found the workflow more effective for covering the “long tail”, such as discovering unseen cases (Q6; $p<0.05$) and defining policy boundaries in realistic scenarios (Q7; $p<0.01$). Participants also emphasized that the workflow clarified their policy decisions where 12 vs. 6 participants rated at 6 or 7 (Q8; $p<0.01$). Furthermore, participants perceived that their adjustments were better reflected by the AI using our workflow (Q9; 13 vs. 6 rated at 6 or 7; $p<0.05$), echoing the stronger alignment performance shown in RQ 4. Lastly, participants were more satisfied overall, with stronger intentions to reuse it for future tasks (Q10; $p<0.01$).
@@ 8. Discussion
Based on the qualitative observations and participant feedback from the user study, we discuss how our interactive system facilitates the iterative co-evolution process and reflect on its limitations, which suggest design implications for future systems.
@@ 8.1.1. Enabling Fine-Grained Specification through Concrete Examples
A significant advantage reported by participants was the ability to define nuanced, domain-specific policies by directly interacting with concrete data examples. This was particularly crucial for ambiguous tasks like content moderation, where abstract guidelines often fail to capture real-world complexities. By reviewing and labeling diverse examples, participants could more easily identify and articulate the subtle boundaries of their criteria. One participant noted how this process helped them refine their standards: “By creating a diverse test set for ‘Incorrect’ items where the boundary was ambiguous, I was able to consider more detailed criteria.”
This hands-on approach, grounded in tangible data, transformed the abstract task of policy-making into a more manageable and effective process. Another participant confirmed this, stating: “Seeing diverse expressions [of criticism] that I could encounter in real life with my own eyes was very helpful in setting more detailed standards for the acceptable level of criticism.” This aligns with Policy Maps (Lam et al., 2025), which showed that surveying concrete cases helps designers define custom regions in the space of model behavior.
@@ 8.1.2. Facilitating Nuanced and Rapid Co-evolution
Participants highly valued the system’s interactivity. The rapid feedback loop allowed them to quickly test hypotheses and iterate on their prompts, making the refinement process more efficient and intuitive. One participant remarked: “It was impressive that it processed feedback in real-time and let me modify the prompt command again.”
The comparison table view was essential for helping users understand the concrete impact of their changes. By visualizing the outputs of the old and new prompt instructions simultaneously, users could easily assess whether their modifications led to the desired improvements. This feature directly supported the goal-oriented nature of their task. “Being able to compare the results of the old prompt and the new prompt at a glance was helpful in writing the prompt’s content (e.g., community guidelines) because I could see what difference the new prompt made.” This underscores the importance of not just enabling rapid iteration but also providing tools that make the outcomes of those iterations transparent and understandable.